Can someone please help me with this problem, I am so confused. Thank you in advance!
Find the slope of the graph of x=-4y+1

Answer by rajagopalan(174)

Find the slope of the graph of x=-4y+1
-4y+1 = x
multiply thro out by -1,
4y-1=-x
4y=-x+1
y=(-x/4)+(1/4)
this is in y = mx+c form
where m=-1/4 and c=1/4
Therfore the slope is -(1/4).
good day.